### v1

# Estimating the Reynolds number dependence of the chaotic attractor dimension in two-dimensional Kolmogorov flow

Abstract
Deep autoencoder neural networks can generate highly accurate, low-order representations of turbulence. We design a new family of autoencoders which are a combination of a ‘dense-block’ encoder-decoder structure (Page et al, J. Fluid Mech. 991, 2024), an ‘implicit rank minimization’ series of linear layers acting on the embeddings (Zeng et al, Mach. Learn. Sci. Tech. 5, 2024) and a full discrete+continuous symmetry reduction. These models are applied to two-dimensional turbulence in Kolmogorov flow for a range of Reynolds numbers $25\leq Re\leq 400$, and used to estimate the dimension of the chaotic attractor, $d_{\mathcal{A}}(Re)$. We find that the dimension scales like $\sim Re^{1/3}$ — much weaker than known bounds on the global attractor which grow like $Re^{4/3}$. In addition, two-dimensional maps of the latent space in our models reveal a rich structure not seen in previous studies, including multiple classes of high-dissipation events at lower $Re$ which guide bursting trajectories. We visualize the embeddings of large numbers of “turbulent” unstable periodic orbits, which the model indicates are distinct (in terms of features) from any flow snapshot in a large turbulent dataset, suggesting their dynamical irrelevance. This is in sharp contrast to their appearance in more traditional low-dimensional projections, in which they appear to lie within the turbulent attractor.
I Introduction
Nonlinear partial differential equations (PDEs), such as the Navier–Stokes equations that govern turbulent fluid motion, are fundamental in modeling complex dynamics and often exhibit spatiotemporal chaos in parameter regimes of practical interest. Although the state spaces of PDEs are formally infinite-dimensional, it is argued that solutions of the Navier-Stokes equations collapse onto a finite-dimensional inertial manifold, denoted here by $\mathcal{M}$, due to the dissipative action of viscosity [1, 2, 3, 4]. Although the existence of inertial manifolds for the Navier-Stokes equations has not been proven [5], approximate inertial manifolds, which attract all orbits within a thin neighborhood of the global attractor, have been constructed in two spatial dimensions [6]. The inertial manifold encapsulates the essential features of the system, with an intrinsic dimension $d_{\mathcal{M}}$, typically much smaller than the dimension $d_{\omega}$ of the state space of the discretized PDE.
[…]
Our focus in this work is on generating low-dimensional embeddings for the chaotic attractor $\mathcal{A}\subset\mathcal{A}_{G}$ in two-dimensional Kolmogorov flow, which describes the long-time evolution of turbulence and which presumably requires far fewer degrees of freedom than $\mathcal{A}_{G}$. To estimate the scaling $d_{\mathcal{A}}:=\dim(\mathcal{A})$ with Reynolds number, $Re$, we construct highly-accurate autoencoder neural networks in an effort to find robust ‘latent’ coordinate systems. Our objectives are twofold: (1) to estimate the $Re$-dependence of $d_{\mathcal{A}}$ and (2) to use the latent coordinate system to map typical dynamical pathways through different vortical ‘events’ and to understand the coverage (or lack of coverage) of these dynamics by large libraries of recently computed unstable periodic orbits [13, 14].
Numerical computations for the chaotic attractor dimension (by computing Lyapunov exponents and invoking the Kaplan-Yorke conjecture [15]) give some indications as to a scaling with $Re$. The Lyapunov dimension of the attractor of Kolmogorov flow was reported to depend strongly on the wavenumber $n$ chosen for the monochromatic forcing profile, and only weakly on $Re\propto G^{1/2}$ [16]. A follow-on study of Kolmogorov flow found that the Lyapunov dimension is bounded above by the number of degrees of freedom contained in scales larger than the forcing scale, $N_{LS}\propto n^{2}$ [17], though this study considered relatively weak turbulence with $16.4\leq Re\leq 46.2$ in our units. A scaling of the Lyapunov dimension $\sim Re^{0.78}(n/k_{0})^{2}$ was also computed numerically for forced two-dimensional turbulence with large scale damping, where $k_{0}=2\pi/L$ [18], which also has a linear dependence on $N_{LS}$.
Related to these direct numerical estimates, there has been a recent focus on using data-driven approaches to estimate the dimensionality of the inertial manifold of simpler chaotic systems. Linear approaches, such as principal component analysis (PCA) and its extensions (e.g., Sparse PCA [19], Bayesian PCA [20]), tend to overestimate the dimension as they are limited to linear subspaces. Nonlinear techniques (e.g. deep autoencoders) address this limitation to more faithfully learn the structure of the inertial manifold. By examining how the model reconstruction error of the data changes with the dimensionality of the embedding space, autoencoder networks have been used to identify the inertial manifold dimension of the Kuramoto–Sivashinsky (KS) equation [21, 22, 23]. However, this approach requires a manual sweep over multiple latent dimensions and there is no apparent drop-off in reconstruction error for more complex systems and higher spatial dimensions [24, 25]. A recent study has proposed a more systematic approach to determine the dimensionality of the inertial manifold of the KS equation, using the so-called Implicit Rank Minimizing Autoencoder with Weight Decay (IRMAE-WD) [26]. The design of this architecture exploits the phenomenon that gradient-based optimization of deep linear neural networks leads to low-rank solutions (implicit rank minimization) [27], while the weight decay refers to an $L_{2}$ regularization on the weights of the network to further drive down the rank of the learned embedding. The same network was applied to both two-dimensional Kolmogorov flow with a forcing wavenumber of $n=2$ at low $Re$ [28], and two-dimensional Rayleigh-Bénard flow [29] to compute estimates for the dimensionality of these flows.
Many canonical fluid flows exhibit continuous and/or discrete symmetries in the governing equations, which must be accounted for in the training of a nonlinear neural network. The simplest way to encourage the model to learn the symmetries of the equations is via data augmentation, where each symmetry is randomly applied to the training data after each epoch [13, 30] to ensure a robust sampling of the symmetry groups. However, this approach does not guarantee equivariance of the model with respect to the symmetries. In this work, to explicitly reduce a continuous or discrete symmetry, each snapshot is mapped a symmetry-reduced subspace of the full state space. This pre-processing of the data restricts the model to learning a representation within this subspace [31, 28]. The first Fourier mode method of slices can be used to explicitly reduce continuous symmetries of dynamical systems [32]. For discrete symmetries, we adapt a reduction technique that defines an indicator function based on Fourier modes to separate the state space of the governing equation into symmetric charts, and maps all solutions into one of these charts, defining a ‘fundamental chart’ [this was applied to the KS system in 31]. This method of ‘symmetry charting’ has been applied to Kolmogorov flow with forcing wavenumber $n=2$ [28], where an indicator function was defined on specific Fourier modes to define eight symmetry charts representing the eight-fold discrete symmetries. Other work has explicitly accommodated various discrete symmetries via the use of many identical (‘Siamese’) networks to identify a fundamental domain [33].
Other studies have also considered the problem of simulating dynamics in the latent space of an autoencoder – motivated by the extreme computational requirements of direct numerical simulation of turbulence at high $Re$. Kolmogorov flow has been a popular testing ground for these efforts to date. For example, convolutional autoencoders have yielded a latent space representation which revealed a clear separation between the high dissipation bursting dynamics and the low dissipation quiescent dynamics, at moderate Reynolds number $Re$ [34, 13]. Autoencoders have also been combined with echo state networks as a reduced order model to accurately predict the flow at different $Re$ [35]. An imperfect reduced order model for Kolmogorov flow has been complemented with a data-driven correction term to accurately predict extreme high dissipation events [36].
In this work, we will combine convolutional ‘dense blocks’ [13], the IRMAE-WD network approach [27, 31] and a full symmetry reduction step via symmetry charting, to estimate the scaling with $Re$ of an upper bound to the chaotic attractor dimension of the $n=4$ Kolmogorov flow. The combination of the IRMAE model with a full symmetry reduction yields a very rich and interpretable learned latent space, which reveals new physical insight to the dynamics of the flow. In section II, we formulate the equations of motion of two-dimensional Kolmogorov flow and describe the data generation and symmetry reduction. In section III, we describe the autoencoder architecture. In section IV, we report the estimated scaling of the chaotic attractor dimension with $Re$. In section V, we investigate the structure and modal decomposition of the learned embedding spaces of the models.
II Flow configuration
II.1 Numerical simulations
We consider two-dimensional turbulence on a doubly periodic domain, driven by a monochromatic body force in the streamwise direction (Kolmogorov flow). The out-of-plane vorticity $\omega=\partial_{x}v-\partial_{y}u$, where the velocity $\bm{u}=(u,v)$, evolves according to
$\partial_{t}\omega+\bm{u}\cdot\bm{\nabla}{\omega}=\frac{1}{Re}\nabla^{2}\omega-n\cos{ny}.$ (1)
In this non-dimensionalization, a length scale is chosen as the inverse of the fundamental wavenumber of the box $1/k^{*}=L^{*}/2\pi$, where $L^{*}=L^{*}_{x}=L^{*}_{y}$ ( i.e. equal aspect ratio – the asterisk denotes dimensional units here). A time scale is chosen as $1/\sqrt{k^{*}\chi^{*}}$, where $\chi^{*}$ is the amplitude of the forcing in the momentum equation. These length and time scales define the Reynolds number for this configuration $Re\coloneq\sqrt{\chi^{*}/k^{*3}}/\nu$, where $\nu$ is the kinematic viscosity. Throughout this work, the forcing wavenumber is set to $n=4$, as has been common in previous studies [37, 34, 14]. Equation (1) with periodic boundary conditions and $n=4$ forcing waves is equivariant under the action of three symmetries:
  1. 1.
    continuous shifts in the streamwise ($x$) direction, $\mathscr{T}_{s}:\omega(x,y)\to\omega(x+s,y)$,
  2. 2.
    discrete shift-reflects by a half-wavelength in $y$, $\mathscr{S}:\omega(x,y)\to-\omega(-x,y+\pi/4)$ and
  3. 3.
    discrete rotations by $\pi$, $\mathscr{R}:\omega(x,y)\to\omega(-x,-y)$,
which we account for explicitly when estimating the dimension of the turbulent attractor.
Key measures of the flow used throughout this study are the total dissipation rate,
$D\coloneq\frac{1}{Re}\langle|\nabla\bm{u}|^{2}\rangle_{V},$ (2)
and the total production rate,
$I\coloneq\langle u\sin(ny)\rangle_{V},$ (3)
where the average over the volume $V=[0,2\pi]^{2}$ is defined as
$\langle\bullet\rangle_{V}\coloneq\frac{1}{(2\pi)^{2}}\iint_{V}\,\bullet\,d^{2}\bm{x}.$ (4)
[…]
We solve equation (1) using the spectral version of the JAX-CFD solver [39, 40]. The velocity field is computed at each time-step via solution of the Poisson equation $\nabla^{2}\psi=-\omega$, where the streamfunction $\psi$ is related to the induced velocity components via $u=\partial_{y}\psi,v=-\partial_{x}\psi$.
II.2 Datasets
For $Re\leq 200$, training data was generated from 400 trajectories, each of length $10^{3}$ advective time units, with an initial transient period of 50 advective time units discarded. At $Re=400$, we instead computed 4000 trajectories each of length $10^{2}$ (to enable parallelization over trajectories and avoid excessive compute time at high resolution). Data was sampled every advective time unit ($\Delta t^{*}=1/\sqrt{k^{*}\chi^{*}}$ in dimensional units), resulting in a dataset of $4\times 10^{5}$ individual snapshots for each $Re$ value considered. A spatial resolution of $128^{2}$ was sufficient for $Re<200$, while trajectories at $Re=200$ and $400$ were computed with $256^{2}$ and $512^{2}$ gridpoints respectively, before being downsampled to $128^{2}$ for input to the autoencoders. We reserved 10% of the training data as a ‘validation’ dataset, which is not used to update the model weights, but instead is used to track convergence metrics throughout training. Test datasets of $5\times 10^{4}$ snapshots for each $Re$ were generated in the same manner.
II.2.1 Resampling
[…]
Figure 1: The distributions of dissipation rate $D$ at $Re=100$, normalized by the laminar dissipation rate $D_{lam}=Re/(2n^{2})$, is shown for the full training dataset (gray) and the resampled training dataset (blue).
II.2.2 Symmetry reduction
As described in §II.1, Kolmogorov flow is equivariant under continuous shifts in the streamwise direction, under discrete shift-reflects by a half-wavelength in $y$, and under rotations by $\pi$. We account for all symmetries explicitly in a pre-processing step before training the neural networks The continuous translational symmetry is reduced by the first Fourier mode method of slicing [32]. In this method a phase $\phi$ is computed such that the $(k_{x},k_{y})=(1,0)$ Fourier mode of the translated state $\mathscr{T}_{-\phi}\omega$ is purely real.
For the discrete symmetries, we adapt the symmetry charting approach introduced in [28] for $n=2$ Kolmogorov flow. This approach partitions the full state space into symmetrically related charts, and assigns each chart a unique indicator $\mathscr{I}$. We select a single chart to which we map all states by combinations of $\mathscr{S}$ and $\mathscr{R}$, which is done by looking at specific terms in the Fourier transform of the vorticity field,
$\omega(x,y)=\sum_{k_{x}}\sum_{k_{y}}\hat{\omega}(k_{x},k_{y})e^{i(k_{x}x+k_{y}y)}.$
The actions of the three symmetry operations on the Fourier modes are
$\mathscr{T}_{s}:\hat{\omega}(k_{x},k_{y})$ $\to\hat{\omega}(k_{x},k_{y})e^{-ik_{x}s},$ (5)
$\mathscr{S}:\hat{\omega}(k_{x},k_{y})$ $\to-\hat{\omega}(-k_{x},k_{y})e^{ik_{y}\pi/n},$ (6)
$\mathscr{R}:\hat{\omega}(k_{x},k_{y})$ $\to\overline{\hat{\omega}}(k_{x},k_{y}),$ (7)
where $\overline{\bullet}$ denotes complex conjugation. It is clear from (5) that the action of $\mathscr{T}_{s}$ only changes modes with $k_{x}\neq 0$. However, $\mathscr{R}$ changes all modes, while $\mathscr{S}$ changes all modes except for modes $(0,(2l+1)n)$, where $l\in\mathbb{Z}$. As $\mathscr{T}_{s}$ does not affect modes with $k_{x}=0$, we will use specific $k_{x}=0$ modes to specify the indicators of the discrete symmetries, and reduce the discrete symmetries before the translational symmetry. Otherwise, if the discrete symmetries are reduced after the translational symmetry, the (1,0) mode is no longer guaranteed to be purely real.
We consider the $(0,1)$ mode to diagnose the shift-reflect symmetry, splitting the complex plane spanned by the real $\hat{\omega}_{R}(0,1)$ and imaginary $\hat{\omega}_{I}(0,1)$ components of this mode into octants, representing the 8 shift-reflected copies generated by (6). We consider the $(0,4)$ mode for the rotational symmetry, splitting this complex plane into upper and lower halves, i.e. defined by the sign of $\hat{\omega}_{I}(0,4)$. Together, this decomposition results in 16 indicators as indicated in figure 2, representing the 16 symmetric copies generated by both (6) and (7). To reduce the discrete symmetries of a state, we select the ‘fundamental’ domain as $\mathscr{I}=1$. The required combination of symmetry operations to map a state from any $\mathscr{I}$ to $\mathscr{I}=1$ is tabulated in table 1.
Generalization of this approach to any number of forcing waves $n$ is relatively straightforward – it requires the partitioning of the complex plane spanned by $\hat{\omega}_{R}(0,1)$ and $\hat{\omega}_{I}(0,1)$ into $n$ equally-sized sectors, along with consideration of the sign of $\hat{\omega}_{I}(0,n)$ to diagnose rotation.
Figure 2: 16 symmetry charts of $n=4$ Kolmogorov flow, for each of the 16 discrete symmetric copies of a state. The $x$- and $y$-axes denote $\hat{\omega}_{R}(0,1)$ and $\hat{\omega}_{I}(0,1)$ respectively. (Left) $\hat{\omega}_{I}(0,4)>0$ and (right) $\hat{\omega}_{I}(0,4)<0$.
$\mathscr{I}$ Action
1 $\mathbb{I}$
2 $\mathscr{S}^{3}$
3 $\mathscr{S}^{6}$
4 $\mathscr{S}$
5 $\mathscr{S}^{4}$
6 $\mathscr{S}^{7}$
7 $\mathscr{S}^{2}$
8 $\mathscr{S}^{5}$
$\mathscr{I}$ Action
9 $\mathscr{S}^{5}\mathscr{R}$
10 $\mathscr{S}^{2}\mathscr{R}$
11 $\mathscr{S}^{7}\mathscr{R}$
12 $\mathscr{S}^{4}\mathscr{R}$
13 $\mathscr{S}\mathscr{R}$
14 $\mathscr{S}^{6}\mathscr{R}$
15 $\mathscr{S}^{3}\mathscr{R}$
16 $\mathscr{R}$
Table 1: Symmetry operations to map all 16 symmetry charts of $n=4$ Kolmogorov flow to the fundamental chart $\mathcal{I}=1$
III Network architecture
We construct low-rank representations of the turbulent attractor in Kolmogorov flow by combining the DenseNet convolutional autoencoder architecture from a previous study [13] (hereafter PHBK) with the implicit rank minimization and weight decay introduced by the IRMAE-WD architecture [26, 28]. Each deep convolutional autoencoder $\mathscr{A}$ seeks to reconstruct its input such that
$\mathscr{A}(\omega)\equiv[\mathscr{D}\circ\mathscr{W}\circ\mathscr{E}](\omega)\approx\omega,$ (8)
where the encoder $\mathscr{E}:\mathbb{R}^{N_{x}\times N_{y}}\to\mathbb{R}^{d_{z}}$ maps the input vorticity snapshot to a low-dimensional representation, $\mathscr{W}:\mathbb{R}^{d_{z}}\to\mathbb{R}^{d_{z}}$ represents a series of $n$ fully-connected, equally-sized linear layers ( pure matrix multiplication) within the embedding space such that $\mathscr{W}=\mathbf{W}_{1}\mathbf{W}_{2}\cdots\mathbf{W}_{n}$ and $\mathbf{W}_{i}\in\mathbb{R}^{d_{z}\times d_{z}}$, and the decoder $\mathscr{D}:\mathbb{R}^{d_{z}}\to\mathbb{R}^{N_{x}\times N_{y}}$ maps the embedding back to a vorticity snapshot.
The structures of the encoder $\mathscr{E}$ and decoder $\mathscr{D}$ considered here are adapted from PHBK, with so-called ‘dense blocks’ [42, 43] instead of single convolutional layers. The dense blocks allow for much richer feature maps at each scale throughout the autoencoder. Full details on the architecture used here are given in appendix A. The key addition to the architecture is the insertion of fully-connected linear layers into the bottleneck of the autoencoder. The implicit rank minimization of these linear layers drives the output of $\mathscr{W}(\mathscr{E}(\omega))$ to be of minimal rank [27]. The densely connected linear layers to the autoencoder bottleneck do not increase the expressivity of the network, as these layers can be expressed as a multiplication by a square matrix. However, the number of parameters increases to $\sim 6.21\times 10^{6}$ trainable parameters due to the insertion of these linear layers, nearly triple the $\sim 2.15\times 10^{6}$ trainable parameters in the comparable models considered in PHBK. The ‘IRMAE’ approach – while fully deterministic – has been shown to result in latent (or embedding) spaces with many favorable properties reminiscent of more complex approaches (e.g. variational autoencoders), including ability to smoothly interpolate between embeddings and to generate new samples [27].
The networks are trained to minimize the standard mean squared error loss function
$\mathscr{L}=\frac{1}{N_{S}}\sum_{j=1}^{N_{S}}\|\mathscr{A}(\omega_{j})-\omega_{j}\|^{2},$ (9)
using an AdamW optimizer [44] with an initial learning rate of $10^{-4}$ and the default weight decay coefficient of 0.004. This optimizer decouples the weight decay from the adaptive gradient updates, as including an explicit $L_{2}$ regularization term on the trainable parameters in (9) is not effective with Adam-style optimizers [45], nor identical to weight decay [44]. The batch size is set to 8 for $Re\leq 40$, 12 for $Re=60$ and 16 for $Re\geq 80$. Setting relatively small batch size increases the total training time, but also increases the total number of gradient-based optimization steps which we found to to be effective in driving down the rank of the learned embedding. The networks are trained with the AdamW optimizer for at least 1000 epochs for each $Re$. Inspired by previous studies implementing a two-optimizer training process [46], the best model to that point is then further tuned using the AdaGrad optimizer [47] for a further 1000 epochs.
Alternative loss functions, for instance asserting that the time evolved reconstructed vorticity fields must match the time evolved ground truth [48], and spectral regularization terms which penalize poor energy spectrum reconstruction [49], were also considered here but did not result in any substantial improvements to the reconstruction error. We also attempted to fit a second-stage autoencoder network, sharing the same latent representation of the data, to the residual of the reconstructed vorticity fields, motivated by the architecture used to perform machine precision reconstruction of a single turbulent snapshot [50]. However, the error did not appear to have any structure to learn and the second-stage autoencoder was unable to reduce the error any further.
IV Dimension estimation
In this section, we first report the reconstruction error of our model across all $Re$ considered. We then present the estimated scaling for (an upper bound on) the dimension of the chaotic attractor with $Re$.
IV.1 Reconstruction error
We report the relative error for each snapshot $\omega_{j}$ in the test datasets,
$\varepsilon_{j}\coloneq\frac{\|\omega_{j}-\mathscr{A}(\omega_{j})\|}{\|\omega_{j}\|},$ (10)
where the norm is defined as $\|\omega_{j}\|\coloneq\sqrt{(1/4\pi^{2})\iint\omega_{j}^{2}d^{2}\mathbf{x}}$. The average test error for all $Re$ examined is reported in figure 3, with error bars defined by one standard deviation. As expected, the error increases as $Re$ is increased, from $\varepsilon\sim 0.5\%$ at $Re=40$, up to $\varepsilon\sim 3.5\%$ at $Re=400$. These models are more than twice as accurate as those described previously in [13].
Figure 3: The average reconstruction error $\varepsilon$ of the model against $Re$. The error bars at each $Re$ are defined by one standard deviation of the reconstruction error.
The role of some of the new features introduced in the data pre-processing/network training approach is assessed in figure 4. The distribution of errors from the best model from the first stage of training with the AdamW optimizer (red) is notably shifted compared with the best model from the second stage of training with the AdaGrad optimizer (blue). Heuristically, this two-optimizer training process first explores the training loss landscape via the inherent momentum and bias correction of AdamW, before AdaGrad’s feature-adapted learning rate allows it to settle more deeply into the best local minimum encountered in the first training stage. To quantify the impact of the full symmetry reduction preprocessing, we train on a dataset containing all symmetries in which data augmentation is applied at each training epoch by a random application of the discrete and continuous symmetry operations. Even with the data augmentation, the reconstruction error of this model (green) is notably larger than the model trained on fully reduced data, as shown in figure 4. We shall also see in §V that explicitly factoring out the symmetries leads to a rich latent-space structure.
Figure 4: The distribution of the reconstruction error $\varepsilon$ at $Re=40$ is shown for the best model trained with data augmentation and no symmetry reduction (green), the best model from the first AdamW optimizer stage on symmetry reduced data (red), the best model from the second Adagrad optimizer stage on symmetry reduced data (blue).
The distribution of the error and correlation with dissipation rate is examined for a sample of the $Re$values considered in figure 5. For the $Re=40$ results, we also compare the performance of identical architectures trained on dissipation-resampled (color) or not resampled (gray) data at $Re=40$. For a fair comparison, a similar number of training snapshots were used in both cases. The resampling strategy considerably improves the error of the high dissipation events, and surprisingly also slightly improves the error of the low dissipation events. The reconstruction error is still slightly correlated with dissipation rate, especially at higher $Re$.
Figure 5: Histograms of the test datasets at (top) $Re=25,40,80$ and (bottom) $Re=100,200,400$ from left to right, comparing the snapshot dissipation rate $D$ normalized by the laminar dissipation rate $D_{lam}=Re/(2n^{2})$ to the reconstruction error $\varepsilon$ of the snapshot. The colored histograms in each panel correspond to models trained with resampled training data, while the gray histogram at $Re=40$ (also on a logarithmic scale) corresponds to the model trained on the same number of non-resampled data.
We report the test-averaged energy spectrum $R(k)$ of the residual fields $r_{j}=\mathcal{A}(\omega_{j})-\omega_{j}$ generated by the model at $Re=40$ and $Re=100$ in figure 6. We have replaced the traditional single output layer in our architecture by a sequence of linear convolutional layers of decreasing kernel sizes from $(16,16)$ to $(6,6)$ (full details in appendix A). This ‘output stream’ of layers was motivated by the empirical observation that the reconstructed energy spectrum with a single output convolutional layer plateaued incorrectly at high wavenumbers. This plateau is evident in figure 6, where the red residual spectrum is generated by an identical fully-trained architecture with a single linear output layer of filter size $(8,8)$ instead of the output stream. Empirically, the reduction in filter size through the stream tends to sequentially refine at finer spatial scales. Comparing the residual spectra of the output with and without the output stream at $Re=40$, we see a notable improvement at the largest spatial scale, and a slight improvement at the very small spatial scales. We found that using any smaller filter sizes resulted in an over-smoothing of the reconstructed vorticity fields.
Figure 6: The residual spectrum $R(k)$ generated by the model shown in blue at (left) $Re=40$ and (right) $Re=100$. The red line indicates the residual spectrum at $Re=40$ of a model trained with a single linear convolutional output layer of filter size $(8,8)$.
IV.2 Attractor dimension
Here we present the main result of this study – the scaling with $Re$ of the estimated dimension of the chaotic attractor. Implicit rank minimization drives the rank of the embedding space to an approximately minimal rank solution , which we use as (an upper bound for) our numerically estimated dimension. To estimate this minimal rank, we perform a singular value decomposition (SVD) on the covariance of the latent data matrix $\mathbf{Z}$ [26]. The SVD of $\mathbf{Z}\mathbf{Z}^{T}=\mathbf{U}\bm{\Sigma}\mathbf{U}^{T}$ computes the singular vectors $\mathbf{u}_{j}\in\mathbb{R}^{d_{z}}$ where $\mathbf{U}=[\mathbf{u}_{0}\dots\mathbf{u}_{d_{z-1}}]$ (the left singular vectors of $\mathbf{Z}$) and singular values $\sigma_{j}\in\mathbb{R}$ where $\bm{\Sigma}=\text{diag}(\sigma_{0},\dots,\sigma_{d_{z-1}})$ (the square of the singular values of $\mathbf{Z}$). The number of non-negligible singular values in this spectrum is the estimate for $d_{\mathcal{A}}$ in the fully symmetry reduced subspace. Each of these singular vectors represents some important structure of the original data in the latent space, which will be investigated in more detail in §V.
Figure 7: Singular values $\sigma_{i}$ in the singular value decompositions of the covariance of the latent data matrix at $Re=$ 25 (blue), 30 (orange), 35 (green), 40 (red), 60 (purple), 80 (pink), 100 (gray), 200 (brown) and 400 (yellow).
The singular values for each $Re$ considered are reported in figure 7. In all cases there is a sharp drop before $\sigma_{j}$ becomes negligibly small at a critical index $j=c$, corresponding to the rank of the learned, fully symmetry-reduced, latent space. To compute the estimate for $d_{\mathcal{A}}$ in full state space, which we denote by $\tilde{d}_{\mathcal{A}}$, we must add the continuous and discrete symmetries back into the estimate, such that $\tilde{d}_{\mathcal{A}}=c+1+16$. We computed two estimates for $\tilde{d}_{\mathcal{A}}$ at each $Re$ by training two randomly initialized autoencoders from scratch. These estimates were found to be robust to the training – they differed only slightly between models at all $Re$. The smallest estimate at each $Re$ is reported in figure 8.
The non-dimensional Grashof number $G$ in the rigorous upper bound [7, 8] of the global attractor dimension is related to the present definition of $Re$, with monochromatic $n=4$ Kolmogorov forcing, by $G=2^{5/2}\pi^{3}Re^{2}$. This upper bound then scales with $Re$ like $bRe^{4/3}(1+\log{(2^{5/2}\pi^{3}Re^{2})})^{1/3}$, and is indicated by the bold line in figure 8 ($b$ is an unspecified constant). This bound and the lower bound in [9] both scale with the number of degrees of freedom in a two-dimensional domain ($Re^{4/3}$ is proportional to the square of the domain size $L$ over the Kraichnan scale) with a logarithmic correction.
In contrast, our data-driven estimate for the dimensionality of the chaotic attractor explored by long turbulent trajectories exhibits a much weaker scaling at high $Re$. While we observe some non-monotonic behavior in figure 8 at lower $Re$, there is a sharp transition beyond $Re=40$ beyond which an approximate scaling $d_{\mathcal{A}}\sim Re^{1/3}$ is observed – identified with the dotted line in the figure.
Figure 8: The estimated upper bound for the dimension of the turbulent attractor $\tilde{d}_{\mathcal{A}}$ against $Re$. The bold line shows the rigorous upper bound scaling $\sim Re^{4/3}(1+\log{(2^{5/2}\pi^{3}Re^{2})})^{1/3}$ and the dotted line shows the scaling $\sim Re^{1/3}$.
One caveat with our estimates is that our models do not perfectly reconstruct their inputs, and this error does increase with $Re$. However, the markedly weaker scaling with $Re$ is consistent with the formation of coherent structures which are associated with a subset of wavenumber pathways in a Fourier representation. We now explore the representation of the turbulence learned in these models by decoding individual components from the SVD.
V Latent space structure
In this section, we investigate the structure of the latent space learned by the model at various values of $Re$. By explicitly factoring out the symmetries of Kolmogorov flow, the model does not have to learn implicit representations of these symmetry groups. This results in a much richer representation of the dynamics in the fundamental symmetry chart, which can be explicitly related to all symmetric copies.
V.1 Modal structures
We first examine the contribution to the vorticity field reconstruction from each of the singular vectors (modes) in the SVD of $\mathbf{Z}\mathbf{Z}^{T}$. To do this, we project low and high dissipation snapshots onto the $\ell=0$ mode and combinations of this mode with higher modes. The projection operator onto mode $\ell$ in the embedding space for some vorticity field $\omega$ is defined by
$\mathcal{P}_{\ell}(\omega)=\langle\mathbf{u}_{\ell},\mathscr{W}(\mathscr{E}(\omega))\rangle,$ (11)
where $\langle\cdot,\cdot\rangle:\mathbb{R}^{d_{z}}\times\mathbb{R}^{d_{z}}\to\mathbb{R}$ denotes the inner product in the embedding space. The projection onto the non-negligible modes $\ell<c$ defines the co-ordinates of $\omega$ in minimal-rank co-ordinates
$\phi(\omega)\coloneq(\mathcal{P}_{0}(\omega),\dots,\mathcal{P}_{c-1}(\omega))\in\mathbb{R}^{c}.$ (12)
Figure 9: The contributions $\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from mode 0 and $\mathcal{D}(\mathcal{P}_{\ell}(\omega)\mathbf{u}_{\ell}+\mathcal{P}_{0}(\omega)\mathbf{u}_{0})-\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from modes $\ell=1,\dots,6$. The top row shows the input vorticity snapshots at (left panels) $Re=40$ with $D/D_{lam}=0.06,0.12,0.18,0.26$ from left to right, (right panels) $Re=100$ with $D/D_{lam}=0.016,0.032,0.047,0.064$ from left to right. Each column shows the contribution from mode $\ell$ to the reconstruction of the input vorticity snapshot. Each snapshot in a row shares the same colorbar to highlight the activations of different modes across dissipation rates.
Figure 10: The contributions $\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from mode 0 and $\mathcal{D}(\mathcal{P}_{\ell}(\omega)\mathbf{u}_{\ell}+\mathcal{P}_{0}(\omega)\mathbf{u}_{0})-\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from modes $\ell=1,\dots,6$. The top row shows the input vorticity snapshots at (left panels) $Re=200$ with $D/D_{lam}=0.0064,0.0088,0.0102,0.0233$ from left to right, (right panels) $Re=400$ with $D/D_{lam}=0.0022,0.003,0.0035,0.0054$ from left to right. Each column shows the contribution from mode $\ell$ to the reconstruction of the input vorticity snapshot. Each snapshot in a row shares the same colorbar to highlight the activations of different modes across dissipation rates.
To visualize the physical structures spanned by each mode, we decode the projections onto each mode back to physical space. We add the projection onto the $\ell=0$ mode when projecting onto $\ell>0$ modes as otherwise the output $\mathcal{D}(\mathcal{P}_{\ell}(\omega)\mathbf{u}_{\ell})$ is not physically realistic [for a similar discussion, see 34, 13]. Then, we subtract $\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from the decoded field, in an effort to isolate and interpret the contribution from the $\ell>0$ mode in physical space.
The contributions $\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from mode 0 and $\mathcal{D}(\mathcal{P}_{\ell}(\omega)\mathbf{u}_{\ell}+\mathcal{P}_{0}(\omega)\mathbf{u}_{0})-\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from modes $\ell=1,\dots,6$ are shown in figure 9 for a selection of snapshots spanning the full range of dissipation rates at $Re=40$ and $Re=100$. The same modes for $Re=200$ and $Re=400$ are reported in figure 10. At all $Re$ considered, the $\ell=0$ mode bears some resemblance to the time-averaged vorticity state in the symmetry-reduced subspace across the full range of dissipation rates. This structure is also visually similar to many low-dissipation exact solutions in this flow (traveling waves and relative periodic orbits, see [37, 51, 34] – some examples at $Re=40$ are also included in an appendix). There is a weaker projection onto this state for the high dissipation snapshots at $Re=40$, whereas the projection at the higher $Re\in\{100,200,400\}$ is approximately equal across all snapshots presented. Similar to previous studies [34, 13], the $\ell=0$ mode is central to the latent representation – the singular value $\sigma_{0}$ is two orders of magnitude larger than $\sigma_{1}$ at $Re=40$. The $\ell=1$ mode also captures similar structures at both $Re=40$ and $Re=100$, and is dominated by paired diagonal elongated vortex cores of opposite circulation. This mode is visually similar to the most unstable eigenfunction of many low-dissipation exact solutions which resemble the mean flow state (see figure 21 in the appendix). Both $\ell=0$ and $\ell=1$ modes appear to be rotationally symmetric.
At the higher $Re=200$ and $Re=400$, the representation changes somewhat – there is less evidence of the Kolmogorov forcing profile in the decodes of the dominant SVD vectors. Instead, many of the modes capture local features in the vorticity, including dipole/tripolar structures in different orientations (in particular note modes beyond $\ell\geq 3$.
One striking aspect of the decodes at all $Re$ is the sparsity in the representation – some modes are not active for the input snapshots. This is suggestive of a strong separation of modes for different dynamical processes, but is also likely due in part to the discrete symmetry reduction. The mapping to a fundamental chart is based on the first Fourier mode in the vertical. This can mean that snapshots on a trajectory can suddenly go through a ‘jump’ to another chart which suddenly shifts the local vortical features when symmetry reduction is performed (the first Fourier mode may not be apparent by eye, particularly in high dissipation snapshots), and the autoencoder therefore must learn to represent small scale vortices in different areas of the symmetry-reduced snapshots. This is perhaps apparent in the $\ell=3$ mode at $Re=400$ (figure 10) with a dipole in the upper corners of the decode, though in general the features across the modes reported in figures 9 and 10 are all physically distinct. In contrast to similar attempts to visualize physical representations of individual latent directions (e.g. in [13]) the decodes are all smooth and physically interpretable, even at high $Re$.
V.2 Low-dimensional visualizations
Figure 11: Two-dimensional visualizations of the co-ordinate vector (12) at $Re=40$ using the t-SNE algorithm [52]. The $x-$ and $y-$axes represent the two degrees of freedom in the t-SNE visualization, which are determined such that the similarity between neighboring mapped data (measured by t-Student distributions) approximates the similarity in the original variables (measured by Gaussian distributions). Each snapshot is colored by a binned mean normalized dissipation rate $D/D_{lam}$ where $D_{lam}=Re/(2n^{2})$ is the laminar dissipation rate, using 25 bins in both $x-$ and $y-$directions. The dotted black lines indicate level curves of the Gaussian kernel density estimates of the representation, such that increasingly bold lines indicate increasing local density. The spanwise vorticity of representative states indicated by the corresponding marker visualize the distinct dynamical processes contained within the high density clusters of the visualization. The symbols and numbers in the high-dissipation regions are discussed later (see also figure 12).
The sparseness of the representation in the latent SVD basis motivates a lower-dimensional visualization. We present two-dimensional representations of the modal projections of the embeddings (12) for each snapshot in the test dataset using the t-SNE algorithm [52] for $Re=40$, $Re=100$ and $Re=400$. When constructing these visualizations, we also include previously computed libraries of unstable periodic orbits [for full details see 14, 13, 38] and 10 short trajectories of 100 advective time units sampled at intervals of 0.1 time units, when generating the t-SNE visualization. These snapshots had to be included when running the t-SNE algorithm as it is a non-parametric self-supervised algorithm, but the UPOs and trajectory representations are only plotted in the low-dimensional visualizations when explicitly stated in the text. The two-dimensional visualization plane is divided into $25\times 25$ bins, and the mean dissipation rate of of the snapshots in each bin is computed. Each point in the representation represents a snapshot in the dataset, and is colored by this local mean dissipation rate. The contours in figures 11 and 14 indicate level curves of the local density of the visualization, computed via Gaussian kernels [53]. The qualitative features (the clusters) of the output are robust to different initializations of the algorithm.
The two-dimensional visualizations in figures 11 and 14 should be compared to those computed previously in PHBK and [34] these were dominated by the learned shift-reflect symmetry, and the data appeared to form a large octagon with little other structure. The full symmetry reduction here, combined with the IRMAE approach, results in many more t-SNE clusters with distinct embedding features, each of which represents a distinctive ‘class’ of vorticity snapshots in physical space.
At $Re=40$, representative flow states are sampled from a selection of low-dissipation t-SNE clusters and visualized in the right panels of figure 11, with snapshots identified on the t-SNE visualization with markers. The clusters all show the characteristic zig-zag “staircase” patterns in the vorticity, and are distinguished by the number and relative locations of dominant vortices. The detached clusters indicated by the light blue circle and salmon diamond are notably apart from the bulk of the visualization. They sit far from the higher dissipation events and contain the lowest dissipation fields with large patches of quiescent fluid. The clear clustering into many distinct classes of events – not visible in previous work [34, 13] – and apparent distinction of different types of high-dissipation dynamics motivates further investigation into the nature of bursting events and routes into/out of this region of state space.
The high-dissipation events (yellow/orange/red markers in figure 11) are found in two (possibly three) distinct clusters: one surrounded by lower-dissipation events, labeled ‘1’ in figure 11, and another which sits apart in the visualization and is more diffuse, labeled ‘2’. This second, higher-dissipation region may itself be made up of two sub-classes, though the amount of data in this region makes this difficult to discern. The bursting class can be associated with a subset of latent SVD modes $\ell\in L_{d}$. This is done in figure 12, where the most significant modes are identified by computing the mean projection onto each $\ell$ within the class, $\langle\mathcal{P}_{\ell}(\omega)\rangle_{i}$ ($i=1$ or 2) (the average was taken over all snapshots satisfying $D/D_{lam}>D_{c}$ within manually defined boundaries on the t-SNE plane). The subset of modes with the largest difference $\Delta\langle\mathcal{P}_{\ell}\rangle\coloneq\langle\mathcal{P}_{\ell}\rangle_{1}-\langle\mathcal{P}_{\ell}\rangle_{2}$ are highlighted in red in figure 12 – this defines the subset $L_{d}$.
The physical structures spanned by these modes are revealed in figure 12, by visualizing the decode of the sum of the projection onto these modes $\mathcal{D}(\sum_{\ell\in L_{d}}\mathcal{P}_{\ell}(\omega_{i})u_{\ell})$ for representative snapshots $\omega_{i}$ from the two bursting classes which were indicated in the earlier t-SNE plot (figure 11). Class 1 (the calmer high-dissipation cluster) has a stronger contribution from the SVD mode 0, which was associated with the time-averaged mean flow (see figure 9). The signature of this large-scale mode is clearly visible in the decodes of the subset $L_{d}$ in figure 12. In contrast, there is a weaker projection onto this mode for class 2, the higher dissipation class, which explains its location away from the bulk of the visualization in figure 11. Class 2 instead has a stronger projection onto more of the modes $\ell\in L_{d}$, and also excites more large $\ell$ modes which describe smaller-scale vorticity features.
Figure 12: The two different bursting event classes identified by the t-SNE visualization at $Re=40$. Snapshots are extracted belonging to the bursting event class $i=1,2$, and the average projection onto each mode $\langle\mathcal{P}_{\ell}\rangle_{i}$ of class $i$ is computed. (Top) The difference $\Delta\langle\mathcal{P}_{\ell}\rangle\coloneq\langle\mathcal{P}_{\ell}\rangle_{1}-\langle\mathcal{P}_{\ell}\rangle_{2}$ is shown for each mode $\ell$. A subset of modes $\ell\in L_{d}$ (red bars) appears to be mostly responsible for delineating the two bursting event classes. (Bottom) Two representative snapshots $\omega$ indicated by fuchsia (top panels) and cyan (bottom panels) markers from the two classes of bursting events are shown. These snapshots correspond to the fuchsia and cyan circles and crosses in the left panel of figure 11. The decodes $\mathcal{D}(\sum_{\ell\in L_{d}}\mathcal{P}_{\ell}(\omega)u_{\ell})$ are shown directly to the right of each representative snapshot $\omega$.
Plotting highly temporally-resolved trajectories in the t-SNE visualizations can be used to identify routes into and out of the high dissipation regions, i.e. ‘bursting’ events. This allows us to probe whether either bursting class is visited in isolation, and where likely bursts originate from in state space. Within the t-SNE dataset are 10 trajectories of length 100 advective units, and from these we extract and visualize all of the subtrajectories which either lead to bursting events from quiescent events, or vice-versa. These subtrajectories are defined as the continuous sequences of snapshots satisfying $D<D_{c}$ which precede or follow the local dissipation maximum, where $D_{c}$ is the same dissipation rate threshold used for the dataset resampling in §II. We also require that the subtrajectories start from or finish at a sufficiently quiescent state satisfying $D<D_{l}$ and that the local dissipation maximum satisfies $D>D_{h}$, where $D_{l},D_{h}=0.125D_{lam},0.175D_{lam}$ for $Re=40$. The dissipation rate and low-dimensional representation of one such bursting subtrajectory is shown in the middle panel of figure 13, where the dissipation rates of the pre-bursting and post-bursting snapshots are indicated by the opaque markers.
Histograms of all pre- and post-bursting events, visualized over the earlier t-SNE map of the state space, are reported in figure 13. There is a clear distinction between pre- and post-bursting events. The plot reveals that that the ‘calmer’ bursting class is not visited in isolation, but is the route of all high-dissipation events back to quiescent dynamics. This is confirmed by inspecting the full trajectories of bursting events on the t-SNE visualization (not shown), which always first hop between the two clusters within the extreme bursting class 2, before moving to the calmer bursting class 1.
The pre-bursting trajectories always visit the t-SNE cluster marked by a red cross in figure 11, indicating that this cluster encapsulates some important pre-bursting vorticity features. To identify the associated distinguishing latent modes, the mean projection onto each mode $\ell$ for the snapshots within this cluster $\langle\mathcal{P}_{\ell}(\omega)\rangle_{b}$ and for all the snapshots in the test dataset sampled from the turbulent attractor $\langle\mathcal{P}_{\ell}(\omega)\rangle_{turb}$ are computed. The differences between these average projections are visualized in the lower panel of figure 13. This computation indicates that the pre-bursting cluster has a weaker projection onto the $\ell=0$ mode representative of typical low-dissipation states. However, a much stronger projection onto the $\ell=1$ mode, which features staggered diagonal stripes of vorticity (see figure 9) similar to unstable directions associated with many exact solutions which resemble the mean flow state (see figure 21), is the distinguishing feature of the pre-bursting dynamical process within this cluster.
We have also confirmed that pre-bursting and post-bursting regions on the visualization are both characterized by small $|\hat{\omega}(1,0)|$, while the bursting regions have a large value for $|\hat{\omega}(0,n)|$ (not shown). This large value in the bursting region is to be expected, as the production rate of Kolmogorov flow can be expressed by $I=-\hat{\omega}_{R}(0,n)/n$. The small value of $|\hat{\omega}(1,0)|$ in the pre-bursting region and the large value of$|\hat{\omega}(0,n)|$ in the bursting region confirms the significant redistribution of energy between these modes before a bursting event, an observation which was also made in the variational approach of [41].
Figure 13: (Top) Two-dimensional visualizations of the co-ordinate vector (12) at $Re=40$ using the t-SNE algorithm [52]. The dotted black lines indicate level curves of the Gaussian kernel density estimates of the representation, with increasingly bold lines indicating increasing local density. The dotted orange line identifies the level curve of $D/D_{lam}=0.15$. The grayscale histogram indicates the distribution of subtrajectories (left) which precede and (right) follow a bursting event, which is defined by a local maximum in the dissipation rate satisfying $D>0.175D_{lam}$. Each subtrajectory is defined by the closest continuous sequence of snapshots satisfying $D<0.15D_{lam}$ to the local maximum, within 10 advective time units. The subtrajectory is also required to start from or finish at a sufficiently quiescent state satisfying $D<0.125D_{lam}$. (Middle) The dissipation of a sample subtrajectory centered on the bursting event is plotted against time in advective time units, colored according to time. The opaque markers denote the pre- and post-bursting subtrajectories extracted, while the transparent markers indicate the neglected bursting event. The same subtrajectories are shown by the connected, similarly colored points in the t-SNE visualization, demonstrating these sample routes (top left) towards and (top right) away from the bursting event. (Bottom) The difference $\Delta\langle\mathcal{P}_{\ell}\rangle\coloneq\langle\mathcal{P}_{\ell}\rangle_{b}-\langle\mathcal{P}_{\ell}\rangle_{turb}$ is shown for each mode $\ell$, where $\langle\mathcal{P}_{\ell}\rangle_{b}$ and $\langle\mathcal{P}_{\ell}\rangle_{turb}$ denote the mean projection of snapshots belonging to the pre-bursting cluster (red cross in figure 11) and all snapshots in the test dataset, respectively.
Figure 14: Two-dimensional visualizations of the co-ordinate vector (12) at (top) $Re=100$ and (bottom) $Re=400$ using the t-SNE algorithm [52], as in figure 11.
At $Re=100$ (top panel of figure 14), many high-dissipation snapshots are spaced across the lower periphery of the two-dimensional visualization. There is also a distinct detached high-dissipation cluster, indicating a distinct class of events which could not be separated cleanly in earlier models [34, 13]. Although the clustering of data is less obvious than at the lower $Re=40$ value, there are still distinct classes of vorticity snapshots that can be identified in this visualization. Some representative snapshots sampled from these clusters are shown in figure 15. This figure indicates that the classes can be delineated based on the appearance of the large scale vortices in the flow – their number, relative location and strength. The sample from the detached high-dissipation cluster (black circle) again indicates an event where the importance of ‘mean flow’ modes (e.g. $\ell=0$ and 1) is diminished.
Figure 15: Representative snapshots from the dense clusters in the t-SNE visualization of the co-ordinate vector (12) at $Re=100$. The position of each snapshot in the visualization is marked by the correspondingly colored marker in figure 14.
Finally, the t-SNE visualization of the most turbulent flow considered at $Re=400$ (bottom panel of figure 14) no longer appears to cluster events based on a dominant dissipation rate. The high dissipation snapshots are distributed fairly uniformly in the latent t-SNE visualization. The exception here are a couple of small pockets of very high dissipation snapshots at the left periphery of the visualization. This behavior is consistent with the prominence of the large vortex pair associated with the condensate, a state which is apparent in most snapshots and hence leads to a more homogeneous latent representation across the dataset. Some representative snapshots of the high density regions of the t-SNE visualization are presented in figure 16, and all show a large scale vortex pair. Distinctions between snapshots are likely based on smaller scale features (e.g. vortex filaments) and the relative location/shape of the dominant vortices.
Figure 16: Representative snapshots from the dense clusters in the t-SNE visualization of the co-ordinate vector (12) at $Re=400$. The position of each snapshot in the visualization is marked by the correspondingly colored marker in figure 14.
V.3 Periodic orbit coverage
Figure 17: Dissipation rate $D$ against production rate $I$ of the UPOs in the libraries of (left) 174 solutions at $Re=40$ and (right) 252 solutions at $Re=100$, both normalized by the laminar dissipation $D_{lam}=Re/(2n^{2})$. The gray background is the probability density function (PDF) computed from a trajectory of $10^{5}$ samples, separated by 1 advective time unit. The contour levels of the PDF are spaced logarithmically. Solutions deemed ‘dynamically relevant’ by the simple dissipation/energy comparison described in the text are colored in light blue, while those note satisfying the condition are colored in gray.
The motion of a turbulent flow within the turbulent attractor is postulated to be shaped and guided by unstable periodic orbits (UPOs) – exact solutions of the Navier-Stokes equations that each encapsulate some closed cycle of physical processes relevant to the turbulence itself [54, 55, 56]. A key motivation for this viewpoint is the possibility to connect dynamical events in the flow to statistical properties of turbulence, e.g. the cascade(s), via periodic orbit theory [57, 58] which provides weights for expansions of the statistics of a chaotic attractor in terms of the statistics of the UPOs.
Large numbers of solutions have been amassed recently for the Kolmogorov flow studied here [14, 13, 38] at $Re\in\{40,100\}$. These solutions are visualized in figure 17 in terms of their production/dissipation rates, overlayed on the turbulent distribution. These low dimensional visualizations, and others like them, would seem to indicate the ‘dynamical relevance’ [59] of the solutions due to their apparent proximity to the turbulence. However, these visualizations are not nuanced enough to identify the range of events contained in the library of UPOs, and may give a misleading impression as to the relevance of a given solution to the true turbulent dynamics. The t-SNE plots constructed above provide a way to assess the degree to which the range of true vortical events in the flow is covered by the library of solutions.
Figure 18: (Top) Two-dimensional visualizations of the co-ordinate vector (12) at $Re=40$ using the t-SNE algorithm [52]. The dotted black lines indicate level curves of the Gaussian kernel density estimates of the representation, such that increasingly bold lines indicate increasing local density. The dotted orange line indicates the level curve of $D/D_{lam}=0.15$. The points indicate snapshots sampled along the periods of the ‘dynamically relevant’ (light blue) and ‘dynamically irrelevant’ (gray) UPOs in (left) the full library of 174 solutions and (right) the short-period orbits with periods $T_{upo}<5$ advective time units. 50 snapshots are sampled for each UPO. (Bottom) Each row shows equally spaced snapshots (without full symmetry preprocessing) along the periods of three periodic orbits with periods $T_{upo}=23.16,18.77$ and $5.06$, indicated in the t-SNE visualization by the red, purple and green points respectively, from top to bottom, where increasing boldness denotes time evolution along the period of the orbit. The crosses in the right top panel indicate the positions of the corresponding visualized snapshots.
Figure 19: (Top) Two-dimensional visualizations of the co-ordinate vector (12) at $Re=100$ using the t-SNE algorithm [52]. The dotted black lines indicate level curves of the Gaussian kernel density estimates of the representation, such that increasingly bold lines indicate increasing local density. The dotted orange line indicates the level curve of $D/D_{lam}=0.045$. The points indicate snapshots sampled along the periods of the ‘dynamically relevant’ (light blue) and ‘dynamically irrelevant’ (gray) UPOs in the full library of 252 solutions. 50 snapshots are sampled for each UPO.
The $Re=40$ UPOs are visualized within the two-dimensional t-SNE visualization in figure 18. The solutions are colored gray or blue according to a simple a priori ‘dynamical relevance’ criteria of whether the time-averaged dissipation rate $\langle D\rangle$ and energy $\langle E\rangle$ lie within the $1^{st}$ and $99^{th}$ quantiles of the corresponding turbulent distributions. While some high-density regions of the turbulence feature snapshots from the UPOs, there are large areas of high turbulent-snapshot density which do not contain any at all. This is surprising given the robustness of the statistical reconstructions possible with this set of solutions observed in [14, 38].
When only the shorter orbits (here $T<5$) are visualized (see top right panel of figure 18) the picture simplifies further. The low-dissipation, short orbits are confined to the edges of the visualization. On the other hand, there are many high-dissipation UPOs which were previously labeled ‘irrelevant’ (based on $E$ and $D$ – gray symbols) that do overlap with the t-SNE embeddings of turbulent snapshots. This indicates that turbulent orbits may visit the neighborhoods of these solutions despite the mismatch in their production/dissipation values. Snapshots from some individual – modest length – UPOs are also included in figure 18 and are identified in the t-SNE visualization. These UPOs are also colored accordingly in the dissipation versus production rate diagram in the left panel of figure 17. The two longest period orbits, with period $T=23.16$ (red) and $T=18.77$ (purple) visit many of the high density clusters throughout the visualization. In contrast, the $T=5.06$ (green) orbit is shorter and is more localized to a high dissipation region of the visualization. This UPO is the only solution in the library which visits the smaller bursting cluster at $Re=40$ discussed in section V.2.
A similar visualization is conducted for the UPOs computed at $Re=100$ in figure 19. While the vast majority of UPOs have production/dissipation values consistent with that of the turbulence (see figure 17), the t-SNE map indicates that nearly all UPOs are determined by the autoencoder to be distinct from the turbulent dataset. This is quite striking given the overlap in various diagnostic statistics reported here and elsewhere [37, 14, 38]. The apparent lack of similarity can be diagnosed by considering the projections $\mathcal{P}_{\ell}(\omega)$ onto individual latent directions.
Figure 20: The joint distribution of the latent projections $\mathcal{P}_{0}$ and $\mathcal{P}_{2}$ of snapshots sampled from the periods of UPOs deemed ‘dynamically relevant’ due to their dissipation and energy distributions (blue) and snapshots sampled from the turbulent attractor (gray). The densities are colored on a log scale.
To investigate the cause for the separation of these UPOs from the turbulent attractor, the mean projection onto each mode $\ell$ was computed for all the snapshots sampled from the ‘dynamically relevant’ (as determined naively from the dissipation and energy PDFs) UPOs at $Re=100$, $\langle\mathcal{P}_{\ell}(\omega)\rangle_{UPO}$, and for all snapshots in the test dataset sampled from the turbulent attractor at $Re=100$, $\langle\mathcal{P}_{\ell}(\omega)\rangle_{turb}$. Differences were found in the distributions for large numbers of the latent projections, but the most significant distinction between the turbulence and the UPOs was found in the leading mode $\ell=0$. Joint distributions for both the UPO data and turbulence are reported in figure 20 (the other mode is selected as $\ell=2$ here). This figure highlights the substantial difference in the $\ell=0$ distributions – the UPO/turbulence data overlap only slightly in the tails of the distributions. Recall that in §V the $\ell=0$ mode was decoded and shown to be visually similar to the symmetry-reduced, time-average flow state in the turbulence. The dramatic reduction of importance of this mode should also be considered in the context of the continuation study reported in [38], which showed that many $Re=100$ solutions appear to ‘move away’ from the turbulent attractor as $Re$ increases. The results reported here would suggest that perhaps they are not inside the turbulent attractor, but just ‘close by’ at $Re=100$ – consistent with the continuation results.
VI Conclusion
In this study we designed deep autoencoders for learning low-dimensional representations of 2D Kolmogorov flow. The models combined (i) full discrete/continuous symmetry reduction, (ii) dense convolutional input/output streams and (iii) an ‘implicit rank minimizing’ (IRMAE [27, 31]) component. The autoencoders were used to estimate the dimensionality of the turbulent attractor in the two-dimensional Navier-Stokes equations – which was obtained from the rank of the inner-most network representation. An approximate scaling of $\tilde{d}_{\mathcal{A}}\sim Re^{1/3}$ was found in the fully turbulent regime – a much weaker scaling than the known upper/lower bounds on the global attractor [7, 8, 9].
The combination of the IRMAE approach with a full symmetry reduction yielded a rich latent-space structure, with which individual latent SVD modes could be decoded into physically interpretable structures at all $Re$ values. Two-dimensional visualizations of large turbulent datasets revealed distinct clusters corresponding to different ‘classes’ of vorticity snapshots. At $Re=40$, multiple types of high-dissipation events were identified. Visualizations of finely resolved bursting trajectories highlighted a common route to a bursting event involving a large projection onto a particular dominant latent mode, and a common route back to quiescent flow through one of the high dissipation clusters.
The low-dimensional visualizations of turbulence were then used to assess the coverage of the turbulent dynamics by large libraries of unstable periodic orbits at both $Re=40$ and $Re=100$. There was a reasonable representation – with notable gaps – at the lower $Re$ value. However, the visualizations at $Re=100$ revealed that the latent representations of nearly all periodic solutions were distinct from the turbulence in a fundamental way. An analysis of individual latent structures indicated a much weaker projection onto a structure associated with the turbulent mean profile. These distinctions are not apparent in simpler low-dimensional projections (e.g. production/dissipation) and indicate that these solutions may not be embedded in the turbulent attractor.
Autoncoders seek to exactly reconstruct their inputs, and hence many of the $O(Re^{1/3})$ dimensions found here may not correspond directly to dynamical variables. For instance, it is known in three-dimensional turbulence that scales smaller than $\sim 5\pi\eta$ (where $\eta$ is the Kolmogorov length) are ‘slaved’ to the larger scales within the turbulent attractor [60, 61]. The equivalent result in two-dimensional turbulence is unknown at present, but having a model learn only the ‘dynamical’ components of vorticity, determined via an appropriate low-pass filter, would presumably simplify the reconstruction problem and allow for a more robust estimate of the $Re$-dependence. We hope to report on this in the near future.
Appendix A Architecture details
Full details of the architecture used in this work are presented in this section. Dense blocks [42, 43] are used here in place of traditional single convolutional layers in both the encoder and decoder, similar to the approach in [13]. The blocks are groups of convolutional layers where the output of each convolutional layer is concatenated with its input. Each dense block consists of three convolutional layers with 32 filters each. Periodic padding is also applied to each image before a convolution, so that the output of a convolutional layer has the same shape as the input. The ‘GELU’ activation function [62] and batch normalization [63] are applied to the output of convolutional layers. Max pooling layers are used to gradually reduce the dimensionality of the image throughout the encoder, while upsampling layers are applied within the decoder to return the original shape at the decoder output.
The encoder architecture can be summarized by the following sequence of operations:
$\omega$ $\to\text{PC}(8\times 8,32)\to\text{DB}(8\times 8)\to\text{MP}(2\times 2)$
$\to\text{PC}(4\times 4,32)\to\text{DB}(4\times 4)\to\text{MP}(2\times 2)$
$\to\text{PC}(4\times 4,32)\to\text{DB}(4\times 4)\to\text{MP}(2\times 2)$
$\to\text{PC}(4\times 4,32)\to\text{DB}(4\times 4)$
$\to\text{PC}(4\times 4,4)\to\text{Flatten}=\mathscr{E}(\omega),$ (13)
where the first term in the brackets denotes the size of the convolutional filters and the second term denotes the number of filters, ‘PC’ stands for a convolutional layer with periodic padding, ‘DB’ for the dense blocks described above and ‘MP’ for a max pooling layer. The input vorticity field is a single-channel image of size $128^{2}$, hence the output of the encoder $\mathscr{E}(\omega)$ is a flattened vector of length $102416^{2}\times 4=$ 1024.
The vector $\mathscr{E}(\omega)$ is then passed through a series of four fully connected (‘FC’) linear layers with equal input/output dimension. These layers do not include ‘bias’ terms and hence are pure matrix multiplications of the upstream input. The architecture of the bottleneck is thus summarized by
$\mathscr{E}(\omega)\to FC(1024)^{4}=\mathscr{W}(\mathscr{E}(\omega)).$ (14)
This output $\mathscr{W}(\mathscr{E}(\omega))$ is the latent representation learned by the autoencoder. The impact of the internal linear operations is to drive the representation to low rank, something which has also been empirically observed in other studies [27, 31].
To decode this latent representation back to physical space, $\mathscr{W}(\mathscr{E}(\omega))$ is reshaped into an ‘image’ with four channels of dimension $16^{2}$. The structure of the decoder used then essentially mirrors the encoder (replacing max pooling layers with upsampling layers), with the exception of the addition of an output stream of linear periodic convolutional layers (‘LPC’):
[…]
$\to\text{PC}(4\times 4,32)\to\text{DB}(4\times 4)\to\text{UP}(2\times 2)$
$\to\text{PC}(4\times 4,32)\to\text{DB}(4\times 4)\to\text{UP}(2\times 2)$
$\to\text{PC}(4\times 4,32)\to\text{DB}(4\times 4)\to\text{UP}(2\times 2)$
$\to\text{PC}(8\times 8,32)\to\text{DB}(8\times 8)$
$\to\text{LPC}(16\times 16,1)\to\text{LPC}(12\times 12,1)$
$\to\text{LPC}(8\times 8,1)\to\text{LPC}(6\times 6,1)$
$=\mathscr{D}(\mathscr{W}(\mathscr{E}(\omega)))=\mathscr{A}(\omega).$ (15)
These linear layers result in a more faithful reproduction of the spectral content of the input snapshots (see §III).
Appendix B Unstable eigenfunctions
Figure 21: Out-of-plane vorticity for three low-dissipation exact solutions at $Re=40$ (left), alongside the real (middle) and imaginary (right) components of the leading unstable eigenfunction. (Top) A traveling wave solution with a wavespeed of 0.0198. (Middle) A UPO with period $T=12.207$ and translational shift of $s=0.243$. (Bottom) A UPO with $T=5.81$ and $s=0.115$.
We include here three examples of low-dissipation exact solutions at $Re=40$, along with a visualization of their leading unstable eigenfunctions. These snapshots are reported in figure 21 and should be compared with decodes of the leading SVD modes reported in figure 9.
Acknowledgements.
This research has been supported by the UK Engineering and Physical Sciences Research Council through the MAC-MIGS Centre for Doctoral Training (EP/S023291/1). Computations were performed on the Cirrus UK National Tier-2 HPC Service at EPCC (http://www.cirrus.ac.uk). JP acknowledges support from UKRI Frontier Guarantee Grant EP/Y004094/1. We are grateful to Masanobu Inubushi for pointing us to reference [9] with a lower bound for the dimension of the global attractor. JP also thanks Rich Kerswell for useful discussions and Mike Graham for helpful discussion on the IRMAE method.